OpenStudy (anonymous):

Express the series below in summation notation for the specified number of terms.

OpenStudy (anonymous):

\[\sum_{n=0}^{10} An\]

OpenStudy (anonymous):

6. 0 + 1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 + 9

OpenStudy (anonymous):

would the sigma be correct for the series

OpenStudy (anonymous):

Remember, your ending point is 10. Also what happened to A?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

it increases by 1

OpenStudy (anonymous):

Well, in the original problem it was being multiplied by your variable n. Because it's a common factor, you can factor it out after you write out the series.

OpenStudy (anonymous):

how would the sigma look like then

OpenStudy (anonymous):

\[\sum_{n=0}^{10}An= An_0 + An_1 + An_2 ... + An_n=A(n_0+n_1+n_2...+n_n)\] Or in this case, another way to look at it is: \[ \sum_{n=0}^{10}An=A \sum_{n=0}^{10}n \]

OpenStudy (anonymous):

That's a slightly generalized expression. It would probably be better to generalize it like: \[ \sum_{n=0}^{i}An =A \sum_{n=0}^{i}n =An_0+An_1+An_2+....+An_i=A(n_0+n_1+n_2+....+n_i)\]